# Experience Replay with Random Reshuffling

### v1

@@ Abstract
Experience replay is a key component in reinforcement learning for stabilizing learning and improving sample efficiency. Its typical implementation samples transitions with replacement from a replay buffer. In contrast, in supervised learning with a fixed dataset, it is a common practice to shuffle the dataset every epoch and consume data sequentially, which is called random reshuffling (RR). RR enjoys theoretically better convergence properties and has been shown to outperform with-replacement sampling empirically. To leverage the benefits of RR in reinforcement learning, we propose sampling methods that extend RR to experience replay, both in uniform and prioritized settings. We evaluate our sampling methods on Atari benchmarks, demonstrating their effectiveness in deep reinforcement learning. footnotetext: Code: https://github.com/pfnet-research/errr.
@@ 1 Introduction
To explore this possibility, we propose two novel methods that integrate RR. For uniform experience replay, our method RR-C (random reshuffling with a circular buffer) applies reshuffling to the indices of a circular buffer so that each index is sampled exactly once per epoch, thereby reducing variance of sample counts and ensuring a balanced use of past experiences. For prioritized experience replay, we introduce RR-M (random reshuffling by masking), which adapts RR to settings with dynamic transition priorities by dynamically masking transitions that have been sampled more often than expected. This strategy effectively approximates an epoch-level without-replacement sampling even as priorities continuously change.
@@
  • For uniform experience replay, we propose RR-C, which applies RR to circular buffer indices rather than transitions themselves, ensuring balanced utilization while maintaining computational efficiency and transition freshness.
@@
  • We empirically demonstrate through experiments on Atari benchmarks that both RR-C and RR-M provide modest but consistent performance improvements across different deep RL algorithms: DQN, C51, and DDQN+LAP.
@@ 3.1 Sampling methods for experience replay
Given the probabilities of transitions, flexibility remains in how minibatches are sampled according to these probabilities. We summarize the sampling methods of experience replay in popular RL code bases in Table 3 in the supplementary materials. The most common method for both uniform and prioritized experience replay is to sample transitions with replacement, where the same transition can be sampled multiple times in a minibatch. The sample code for uniform experience replay is shown as sample_with_replacement in Figure 1. Other approaches include within-minibatch without-replacement sampling (shown as sample_without_replacement in Figure 1) and stratified sampling that segments the range of query values into strata and samples from each stratum, both of which prevent sampling the same transition multiple times in a minibatch.
@@ 3.2 Uniform experience replay with random reshuffling
  • Equivalence to RR: If the transitions in the buffer are fixed, it should work just like RR in SL.
  • Freshness: Once a new transition is added to the buffer, it should be ready to be sampled in the next minibatch.
@@
The sample code of RR-C is shown as sample_rrc of Figure 1. Specifically, we create a shuffled list of indices $[0,1,\dots,C-1]$ (rr_buffer in the figure), where $C$ is the capacity of the circular buffer, and sequentially consume the list when we sample a minibatch during an epoch. When we reach the end of the list, we create a new shuffled list of indices and start a new epoch. It is easy to see that each index is sampled exactly once in an epoch, which is the key property of RR. Although old transitions are overwritten by new transitions in a circular buffer as RL training proceeds, if a transition stays in the buffer for at least $n$ whole epochs, we can guarantee that it is sampled at least $n$ times. If a sampled index has not been assigned to a transition yet, we simply skip it and continue with the next.
@@ 3.3 Prioritized experience replay with random reshuffling
  • Equivalence to RR with repeated data: If the priorities of transitions are all rational numbers, we can consider an epoch where each transition appears exactly how many times it should be sampled in expectation. For example, given three transitions whose priorities are fixed to [1, 0.5, 2], for an epoch of length 7, the sampling method should sample the first transition twice, the second once, and the third four times.
  • Freshness: Same as in uniform experience replay.
@@
We propose a method that satisfies the above properties by masking the priorities of transitions that have been oversampled relative to expectation, which we call RR by masking (RR-M). The sample code of this method is shown as sample_rrm in Figure 2. It keeps actual and expected sample counts of transitions and updates them every time a minibatch is sampled. When a transition is deemed oversampled, it is masked by setting its priority to a tiny value [^2] In our experiments, we multiply the original priority by 1e-8, not 0, so that the algorithm would work if all the transitions were deemed oversampled due to numerical errors. until its expected count catches up. When old transitions are overwritten by new transitions, we reset the actual and expected counts of the overwritten transitions to zero and scale the expected counts so their sum matches the sum of actual counts. When we sample a minibatch, we do within-minibatch without-replacement sampling according to the masked priorities. This can be achieved by querying the sum tree in a batch manner, removing the duplicated indices if any, temporarily masking the priorities of the sampled transitions, resample only the amount that is less than the desired minibatch size, and repeating the process until the minibatch is filled. We include example trajectories of the behavior of RR-M in Table 4 in the supplementary materials.
This method is also easy to implement and has no hyperparameters to be tuned, but it is relatively computationally expensive because every time we sample a minibatch, we need to update the counts of transitions, which takes $O(n)$ time, where $n$ is the number of transitions in the replay buffer. In practice, it is easy to parallelize the computation of the counts and oversampled indices, e.g., with GPUs. Having two separate sum trees to keep both correct priorities and masked priorities also helps minimize the computational overhead as it has to update a mask of a transition only when it changes whether it is oversampled. We discuss the computational cost in more detail in Section C in the supplementary materials.
@@
Figure 2: A Python code example of our sampling method for prioritized experience replay, RR-M. This simplified code is for illustrative purposes and is not identical to actual implementations.
@@ 4.1 Experience replay simulations
Similarly, the right side of Figure 3 shows the distributions of sample counts in prioritized experience replay simulations. We simulate non-uniform and dynamic priorities; the priority of each transition is set to $p_{t}=(t\bmod 25)+5$ when added to the buffer and decays by a factor of 0.8 every time it is sampled, where these values are chosen arbitrarily for illustration. Again, we observe that RR-M realizes lower-variance sample counts without significant bias. We include additional simulation results with different set of parameters in Section D.1 in the supplementary materials.
Figure 3: Distributions of sample counts in experience replay simulations. We ran 100-timestep simulations with different random seeds for each configuration. For each transition, we visualize how many times it is sampled during a simulation. Solid lines represent mean sample counts, thick shaded areas represent mean±$\pm$±stdev, and light shaded areas represent minimum to maximum over 1000 simulations. WR: with-replacement sampling, WOR: without-replacement sampling, RR-C: RR with a circular buffer, RR-M: RR by masking.
From these simulations, we observe that both our RR-based sampling methods can reduce the variance of sample counts in experience replay, which is expected to enable lower-variance gradient estimation and improve learning stability. We verify the effectiveness in deep RL settings in the following experiments.
@@ 4.2 Deep RL with uniform experience replay
We run the DQN (Mnih et al., 2015) and C51 (Bellemare et al., 2017) algorithms and compare the performance of different sampling methods. We base our implementation on the CleanRL library (Huang et al., 2022) and keep the hyperparameters the same as the default values. We run each experiment with 10 random seeds and report the final performance, measured by the mean score over the last 100 episodes. Learning curves are shown in the supplementary materials.
Table 1: Final performance of C51 and DQN with with-replacement sampling (WR) and RR-C. Final performance is measured by the mean episodic returns of the last 100 training episodes and reported as the mean over 10 random seeds with higher values bold. The p-values are computed by Welch’s t-test. *: $p<0.05$, **: $p<0.01$.
@@
Table 1 shows the final performance of C51 and DQN with different sampling methods. Both C51 and DQN with RR-C outperform their with-replacement sampling counterparts in the majority of games, which demonstrates the effectiveness of RR in experience replay. The effect size of RR is modest in most games, but it is consistent across different games and algorithms.
@@
We also evaluate the performance of C51 with within-minibatch without-replacement sampling and observe that it is no better than with-replacement sampling, which suggests that the benefits of RR come from the buffer-level without-replacement sampling. This is not surprising because sampling with replacement in this setting rarely samples duplicate transitions in the same minibatch; the capacity of the buffer is set to 1 million, which is much larger than the minibatch size of 32. The final performance is shown in Figure 7 in the supplementary materials.
@@ 4.3 Deep RL with prioritized experience replay
Table 2 shows the final performance of DDQN+LAP with different sampling methods. The left side of the table compares with-replacement sampling and RR-M. Although the differences are less pronounced compared to uniform experience replay experiments, RR-M outperforms with-replacement sampling 8 out of 10 games, which suggests that RR can also be beneficial in prioritized experience replay.
@@
Stratified sampling is sometimes used in prioritized experience replay, querying a different segment of the sum tree for each transition in a minibatch. Since priorities are stored in a sum tree in chronological order, stratified sampling ensures diversity in age of transitions in a minibatch, which could have a different benefit for learning. The right side of Table 2 compares stratified sampling and RR-M+ST, where stratified sampling is used instead of within-minibatch without-replacement sampling. We observe that RR-M+ST outperforms stratified sampling in 8 out of 10 games, which suggests that RR can be beneficial even when stratified sampling is used.
Table 2: Final performance of DDQN+LAP with with-replacement sampling (WR), RR-M, stratified sampling (ST), and RR-M+ST (ST applied to RR-M). Format follows Table 1.
@@ 5 Related work
Closely related to our work are combined experience replay (CER) (Zhang & Sutton, 2017) and corrected uniform experience replay (CUER) (Yenicesu et al., 2024). CER ensures the most recent transitions are always included in a minibatch, aiming to mitigate the negative effects of a large replay buffer. CUER is a variant of prioritized experience replay that prioritizes transitions that have been sampled less often by reducing the priority every time a transition is sampled. Similarly to our work, both CER and CUER could reduce the variance in the number of times each transition is sampled, as CER ensures that each transition is sampled at least once and CUER reduces the priority of transitions that have been sampled more often than others. However, our RR-based sampling methods aim to eliminate such unnecessary variance by either enforcing that each transition is sampled exactly once per epoch (RR-C) or detecting and excluding the oversampled transitions (RR-M). While CER and CUER favor recent transitions by design, our RR-based sampling methods do not have such a bias and can be applied to both uniform and prioritized experience replay.
@@ 6 Conclusion
In this work, we introduced two sampling methods that extend random reshuffling to uniform and prioritized experience replay, respectively. Our empirical results on Atari benchmarks demonstrate that introducing RR to experience replay yields modest but consistent improvements across different RL algorithms. Importantly, our RR-based sampling methods are straightforward to implement and can safely replace with-replacement sampling in existing RL code bases, making them practical for RL practitioners.
@@ Appendix A Sampling methods of experience replay in popular RL code bases
We summarize the sampling methods of experience replay in popular RL code bases in Table 3.
@@
ClearnRL (Huang et al., 2022) WR [^4] CleanRL depends on Stable-Baselines3’s ReplayBuffer N/A

### v2

# theorem[name=Theorem]

@@ Abstract
Experience replay is a key component in reinforcement learning for stabilizing learning and improving sample efficiency. Its typical implementation samples transitions with replacement from a replay buffer. In contrast, in supervised learning with a fixed dataset, it is a common practice to shuffle the dataset every epoch and consume data sequentially, which is called random reshuffling (RR). RR enjoys theoretically better convergence properties and has been shown to outperform with-replacement sampling empirically. To leverage the benefits of RR in reinforcement learning, we propose sampling methods that extend RR to experience replay, both in uniform and prioritized settings, and analyze their properties via theoretical analysis and simulations. We evaluate our sampling methods on Atari benchmarks, demonstrating their effectiveness in deep reinforcement learning. Code is available at https://github.com/pfnet-research/errr.
@@ 1 Introduction
To explore this possibility, we propose two novel methods that integrate RR. For uniform experience replay, our method RR-C (random reshuffling with a circular buffer) applies reshuffling to the indices of a circular buffer so that each index is sampled exactly once per epoch, thereby reducing the variance of sample counts and ensuring a balanced use of past experiences. For prioritized experience replay, we introduce RR-M (random reshuffling by masking), which adapts RR to settings with dynamic transition priorities by dynamically masking transitions that have been sampled more often than expected. This strategy effectively approximates an epoch-level without-replacement sampling even as priorities continuously change.
@@
  • For uniform experience replay, we propose RR-C, which applies RR to circular buffer indices rather than transitions themselves, ensuring balanced utilization while maintaining computational efficiency.
@@
  • We analyze our methods both theoretically and through simulations, illustrating their bias and variance properties compared to standard methods.
  • We empirically demonstrate through experiments on Atari benchmarks that both RR-C and RR-M provide modest performance improvements across different deep RL algorithms: DQN, C51, and DDQN+LAP.
@@ 3.1 Sampling methods for experience replay
Given the probabilities of transitions, flexibility remains in how minibatches are sampled according to these probabilities. We summarize the sampling methods of experience replay in popular RL code bases in table˜3. The most common method for both uniform and prioritized experience replay is to sample transitions with replacement, where the same transition can be sampled multiple times in a minibatch. The sample code for uniform experience replay is shown as sample_with_replacement in figure˜1. Other approaches include within-minibatch without-replacement sampling (shown as sample_without_replacement in figure˜1) and stratified sampling that segments the range of query values into strata and samples from each stratum, both of which prevent sampling the same transition multiple times in a minibatch.
@@ 3.2 Uniform experience replay with random reshuffling
  • Equivalence to RR: If the transitions in the buffer are fixed, the sampling method should work just like RR in SL.
  • Variance reduction: Even if the transitions in the buffer are changing, the sampling method should reduce the variance of how many times each transition is sampled while not introducing significant bias.
@@
The sample code of RR-C is shown as sample_rrc of figure˜1. Specifically, we create a shuffled list of indices $0,1,\dots,C-1]$ (rr_buffer in the figure), where $C$ is the capacity of the circular buffer, and sequentially consume the list when we sample a minibatch during an epoch. When we reach the end of the list, we create a new shuffled list of indices and start a new epoch. It is easy to see that each index is sampled exactly once in an epoch, which is the key property of RR. Although old transitions are overwritten by new transitions in a circular buffer as RL training proceeds, if a transition stays in the buffer for at least $n$ whole epochs, we can guarantee that it is sampled at least $n$ times. If a sampled index has not been assigned to a transition yet, we simply skip it and continue with the next.
@@ 3.3 Prioritized experience replay with random reshuffling
  • Epoch equivalence to RR with repeated data: If the priorities of transitions are all rational numbers, we can consider an epoch where each transition appears exactly how many times it should be sampled in expectation. For example, given three transitions whose priorities are fixed to [1, 0.5, 2], for an epoch of length 7, the sampling method should sample the first transition twice, the second once, and the third four times.
  • Variance reduction: Similarly to the uniform case, even if the transitions in the buffer are changing and their priorities are non-stationary, the sampling method should reduce the variance of how many times each transition is sampled while not introducing significant bias.
@@
We propose a method that satisfies the above properties by masking the priorities of transitions that have been oversampled relative to expectation, which we call RR by masking (RR-M). The sample code of this method is shown as sample_rrm in figure˜2. It keeps actual and expected sample counts of transitions and updates them every time a minibatch is sampled. When a transition is deemed oversampled, it is masked by setting its priority to a tiny value [^2] In our experiments, we multiply the original priority by 1e-8, not 0, so that the algorithm would work if all the transitions were deemed oversampled due to numerical errors. until its expected count catches up. When old transitions are overwritten by new transitions, we reset the actual and expected counts of the overwritten transitions to zero and scale the expected counts so their sum matches the sum of actual counts. When we sample a minibatch, we do within-minibatch without-replacement sampling according to the masked priorities. This can be achieved by querying the sum tree in a batch manner, removing the duplicated indices if any, temporarily masking the priorities of the sampled transitions, resample only the amount that is less than the desired minibatch size, and repeating the process until the minibatch is filled. We include example trajectories of the behavior of RR-M in table˜4.
This method is also easy to implement and has no hyperparameters to be tuned, but it is relatively computationally expensive because every time we sample a minibatch, we need to update the counts of transitions, which takes $O(n)$ time, where $n$ is the number of transitions in the replay buffer. In practice, it is easy to parallelize the computation of the counts and oversampled indices, e.g., with GPUs. We discuss the computational cost in more detail in appendix˜D.
@@
Figure 2: A Python code example of our sampling method for prioritized experience replay, $R$-M. This simplified code is for illustrative purposes and is not identical to actual implementations.
Our theoretical analysis shows that RR-M is biased even for later timesteps. Yet, under a certain simplified setting, it can be proved that its relative error to the expected sample counts vanishes and achieves lower variance asymptotically. See appendix˜B for details of the analysis and proofs.
@@ 4.1 Experience replay simulations
Similarly, the right side of figure 3 shows the distributions of sample counts in prioritized experience replay simulations. We simulate non-uniform and dynamic priorities; the priority of each transition is set to $p_{t}=(t\bmod 25)+5$ when added to the buffer and decays by a factor of 0.8 every time it is sampled, where these values are chosen arbitrarily for illustration. Again, we observe that RR-M realizes lower-variance sample counts without significant bias. We include additional simulation results with different set of parameters in section˜E.1.
[…]
From these simulations, we observe that both our RR-based sampling methods can reduce the variance of sample counts in experience replay without introducing significant bias, which complements our theoretical analysis.
@@ 4.2 Deep RL with uniform experience replay
We run the DQN (Mnih et al., 2015) and C51 (Bellemare et al., 2017) algorithms and compare the performance of different sampling methods. We base our implementation on the CleanRL library (Huang et al., 2022) and keep the hyperparameters the same as the default values. We run each experiment with 10 random seeds and report the final performance, measured by the mean score over the last 100 episodes. Learning curves are shown in section˜E.4.
Table 1: Final performance of C51 and DQN with with-replacement sampling (WR) and $R$-C. Final performance is measured by the mean episodic returns of the last 100 training episodes and reported as the mean over 10 random seeds with higher values bold. The p-values are computed by Welch’s t-test. *: $p<0.05$, **: $p<0.01$.
@@
Table˜1 shows the final performance of C51 and DQN with different sampling methods. Both C51 and DQN with RR-C outperform their with-replacement sampling counterparts in the majority of games, which demonstrates the effectiveness of RR in experience replay.
@@
We also evaluate the performance of C51 with within-minibatch without-replacement sampling and observe that it is no better than with-replacement sampling, which suggests that the benefits of RR come from the buffer-level without-replacement sampling. This is not surprising because sampling with replacement in this setting rarely samples duplicate transitions in the same minibatch; the capacity of the buffer is set to 1 million, which is much larger than the minibatch size of 32. The final performance is shown in figure 7.
@@ 4.3 Deep RL with prioritized experience replay
Table˜2 shows the final performance of DDQN+LAP with different sampling methods. The left side of the table compares with-replacement sampling and RR-M. Although the differences are less pronounced compared to uniform experience replay experiments, RR-M outperforms with-replacement sampling in 8 out of 10 games, which suggests that RR can also be beneficial in prioritized experience replay.
@@
Stratified sampling is sometimes used in prioritized experience replay, querying a different segment of the sum tree for each transition in a minibatch. Since priorities are stored in a sum tree in chronological order, stratified sampling ensures diversity in age of transitions in a minibatch, which could have a different benefit for learning. The right side of table˜2 compares stratified sampling and RR-M+ST, where stratified sampling is used instead of within-minibatch without-replacement sampling. We observe that RR-M+ST outperforms stratified sampling in 8 out of 10 games, which suggests that RR can be beneficial even when stratified sampling is used.
Table 2: Final performance of DDQN+LAP with with-replacement sampling (WR), $R$-M, stratified sampling (ST), and $R$-M+ST (ST applied to $R$-M). Format follows table˜1.
@@ 5 Related work
Several studies have investigated sampling methods for experience replay. Panahi et al. (2024) reported within-minibatch without-replacement sampling has a minor benefit over sampling with replacement in prioritized experience replay in tabular settings with smaller buffer sizes. Closely related to our work are combined experience replay (CER) (Zhang & Sutton, 2017) and corrected uniform experience replay (CUER) (Yenicesu et al., 2024); CER ensures the most recent transitions are always included in a minibatch to mitigate the negative effects of a large replay buffer, while CUER, a variant of prioritized experience replay, prioritizes transitions that have been sampled less often by reducing the priority every time a transition is sampled. Similarly to our work, both CER and CUER could reduce the variance in the number of times each transition is sampled, as CER ensures that each transition is sampled at least once and CUER reduces the priority of transitions that have been sampled more often than others. However, our RR-based sampling methods aim to eliminate such unnecessary variance by either enforcing that each transition is sampled exactly once per epoch (RR-C) or detecting and excluding the oversampled transitions (RR-M). While CER and CUER favor recent transitions by design, our RR-based sampling methods do not have such a bias and can be applied to both uniform and prioritized experience replay.
@@ 6 Conclusion
In this work, we introduced two sampling methods that extend random reshuffling to uniform and prioritized experience replay, respectively. Our theoretical analysis and simple simulations suggest that our methods may reduce the variance in the number of times each transition is sampled without introducing significant bias. Our empirical results on Atari benchmarks demonstrate that introducing RR to experience replay yields modest improvements across different RL algorithms. Importantly, our RR-based sampling methods are straightforward to implement and can safely replace with-replacement sampling in existing RL code bases, making them practical for RL practitioners.
@@ Appendix A Sampling methods of experience replay in popular RL code bases
We summarize the sampling methods of experience replay in popular RL code bases in table˜3.
@@
CleanRL (Huang et al., 2022) WR [^4] CleanRL depends on Stable-Baselines3’s ReplayBuffer N/A